Nikki drops her doll from the rooftop 10 m above the ground. Neglecting air resistance, how long it will take to reach the ground?


1
Expert's answer
2021-11-23T10:31:07-0500
"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\times10\ m}{9.8\ \dfrac{m}{s^2}}}=1.43\ s."